OpenStudy (anonymous):

How many solutions does the system have? x= -4y + 4 and 2x + 8y = 8

OpenStudy (shamil98):

substitute for x in the second equation with the given value of x and solve for y. x = -4y + 4 2(-4y+4) + 8y = 8

OpenStudy (anonymous):

Thank you, I tried but it looked to me that there were only two ways to solve this.

OpenStudy (shamil98):

What did you get?

OpenStudy (shamil98):

when everything on both sides equal the same thing like 2x + 5=2x + 5 or 3x - 5 = 3x - 5 you have infinite solutions.

OpenStudy (shamil98):

in this case: 2(-4y+4) + 8y = 8 distribute: -8y + 8 + 8y = 8 simplify 8 = 8 0=0 you have infinite solutions.

OpenStudy (anonymous):

I think I miss one of the steps. Okay. I forgot to simplify

OpenStudy (anonymous):

Thank you for helping understand this better.